40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 19 19:18:53 2017, Kaustubh, Summary, General, Testing of the new Photo Detectors ET-3010 and ET-3040 IMG_20170519_173247922.jpgIMG_20170519_173253252.jpgIMG_20170519_173300174.jpgPD_test_setup.png
    Reply  Mon May 22 18:20:27 2017, Kaustubh, Summary, General, Testing of the new Photo Detectors ET-3010 and ET-3040 ET-3040_test.zipET-3040_test.pdf
Message ID: 12999     Entry time: Fri May 19 19:18:53 2017     Reply to this: 13005
Author: Kaustubh 
Type: Summary 
Category: General 
Subject: Testing of the new Photo Detectors ET-3010 and ET-3040 

Motivation:

I got some hands-on-experience on using RF photodetectors and the Network Analyzer from Koji. There were newly purchased RF photodetectors from Electro-Optics Technology, Inc.. These were InGaAs Photodetectors with model no.: 120-10050-0001(ET-3010) and 120-10056-0001(ET-3040). The User Guide for the two detectors can be found here. This is the first time we bought the ET-3010 model PD for the 40m lab. It has an operation bandwith >1.5GHz(not tested yet), much higher than other PDs of its kind. This can be used for detecting the output as we 'sweep' the laser frequency for getting data on the optical cavities and the resonating modes inside the cavity. We just tested out the ET-3040 model today but will test out the ET-3010 next week.

Tools and Machines Used:

We worked on the optical bench right in front of the main entrance to the lab. We put the cables, power chords, etc. to their respective places. We used screws, poles, T's, I's, multimeter, Network/Spectrum Analyzer(along with the moving table), a lab computer, Oscilloscope, power supply and the aforementioned PDs for our testing. We took these items from the stack of tools at the Y-arm and the boxes of various different labelled palced near the X-arm. We moved the Network Analyzer(along with the bench) from near the Y-arm to our workplace.

Procedure:

I will include a rough schematic of the setup later.

We alligned the reference PD(High Speed Photoreceiver model 1611) and the test PD(ET-3040 in this case) to get optimal power output. We had set the pump current for the laser at 19.5mA which produced a power of 1.00mW at the output of the fiber couple. At the reference detector the measured voltage was about 1.8V and at the DUT it was about 15mV. The DC transimpedance for the reference detector is 10kOhm and its responsivity to 1064 nm is around 0.75A/W. Using this we calculate the power at the reference detector to be 0.24mW. The DC transimpedance for the DUT is 50Ohm and the responsivity of about 0.9A/W. This amounts to a power of about 0.33mW. After measuring the DC voltages, we connected the laser input to the Network Analyzer and gave in an RF signal with -10dBm and frequency modulation from 100 kHz to 500 MHz. The RF output from the Analyzer is coupled to the Reference Channel(CHR) of the analyzer via a 20dB directional coupler. The AC output of the reference detector is given at Channel A(CHA) and the output from the DUT is given to Channel B(CHB). We got plots of the ratios between the reference detector, DUT and the coupled refernce for the Transfer Function and the Phase. We found that the cut-off frequency for the ET3040 model was at arounf 55 MHz(stated as >50MHz in the data sheet). We have stored the data using the lab PC in the directory .../scripts/general/netgpibdata/data.

Result:

The bandwidth of the ET-3040 PD is as stated in the data sheet, >50 MHz.

Precaution:

These PDs have an internal power supply of 3V for ET-3040 and 6V for ET-3010. Do not leave these connected to any instruments after the experiments have been performed or else the batteries will get drained if there is any photocurrent on the PDs.

To Do:

A similar procedure has to be followed in order to test the ET-3010 PD. I will be doing this tentatively on Monday.

Attachment 1: IMG_20170519_173247922.jpg  4.030 MB  | Hide | Hide all
IMG_20170519_173247922.jpg
Attachment 2: IMG_20170519_173253252.jpg  4.016 MB  | Hide | Hide all
IMG_20170519_173253252.jpg
Attachment 3: IMG_20170519_173300174.jpg  3.853 MB  | Hide | Hide all
IMG_20170519_173300174.jpg
Attachment 4: PD_test_setup.png  41 kB  Uploaded Sat May 20 13:55:59 2017  | Hide | Hide all
PD_test_setup.png
ELOG V3.1.3-